OpenStudy (anonymous):

An amusement park earned $1382.50 in the first five minutes of opening; a total of 287 tickets. If each adult ticket was $8 and each student ticket was $3.50, how many adult tickets were sold? How many student tickets?

OpenStudy (anonymous):

I haven't been able to solve this problem for weeks!! Please help!!

OpenStudy (anonymous):

I think it has to do with Matrixies but i cannot figure it out!

OpenStudy (anonymous):

Let adult tickets be a Let student tickets be b a + b = 287 8a + 3.5b = 1382.5

OpenStudy (anonymous):

b = 287 -a substitute this into the second equation

OpenStudy (anonymous):

ok i am trying to work it out...will you please wait to see if i get the right answer?

OpenStudy (anonymous):

84 tickets for adults 203 tickets for kids